Status Report

NASA Space Station On-Orbit Status 19 November 2005

By SpaceRef Editor
November 20, 2005
Filed under , , ,
NASA Space Station On-Orbit Status 19 November 2005

SpaceRef note: This NASA Headquarters internal status report, as presented here, contains additional, original material produced by (copyright © 2005) to enhance access to related status reports and NASA activities.

All ISS systems continue to function nominally, except those noted previously or below.  Saturday — off-duty day for Bill McArthur & Valery Tokarev, except for housekeeping and voluntary work.   

Back on regular wake/sleep cycle after the successful Soyuz relocation yesterday, the crew made final onboard systems restorations for normal crewed operation.


Processing Status
Daily Mission
Return to Flight
Weekly Status
Weekly Science
Daily On-Orbit Status
Daily Crew Timeline
Soyuz | Progress
ISS News | ATV

In the Russian segment (RS), Valery Tokarev reactivated the TEKh-25 payload hardware, consisting of the Skorpio and Spika-S experiments.   [Skorpio’s objective is to monitor environmental radiation parameters with dosimeters inside station compartments at various places and to characterize environmental conditions for conducting scientific and technical experiments.  To operate, Skorpio requires about 6 W of energy.  The Spika-S payload investigates the influence of the space environment, primarily radiation, on selected commercial electronic “next-generation” components. Early Spika experiments on the Mir space station go back as far as 1992.  Its electronics unit, located in the Service Module (SM) working compartment (panel 417), uses about 7 W power, running either in Acquisition or Monitoring mode, and storing its data on a PCMCIA (portable computer memory card international adapter) card.]

In the US segment (USOS), Bill McArthur worked on the computer systems and reconfigured the station’s internal OpsLAN (Operations Local Area Network).

McArthur also worked on updating the SSC (Station Support Computer) A31p Client laptops with the Service Pack 1 software load, required to run some applications when used in a configuration connected to the PCS (Portable Computer System) A31p machines.

All U.S. equipment that had been temporarily moved to the RS was returned by the CDR to regular USOS locations.

With the Elektron deactivated, the FE serviced the Russian BMP harmful impurities removal system, starting the “bake-out” cycle to vacuum on absorbent bed #1 of the regenerable dual-channel filtration system.  Before sleep time today, the bake-out will be terminated.   [Regeneration of each of the two cartridges takes about 12 hours and is conducted only during crew awake periods.  The BMP currently still uses the same vacuum vent valve for regeneration as the Elektron (the latter for venting hydrogen).]

The CDR performed the regular bi-monthly restart of the OCA comm router SSC (Station Support Computer) laptop. 

Bill also collected the periodic reading of the cabin air’s current CO2 concentrations in the SM and Lab, using the U.S. CDMK (Carbon Dioxide Monitoring Kit, #1015), for calldown, along with its battery status, taken after pump start-up, for use in trending analyses.

Tokarev worked with TsUP/Moscow specialists on a checkout of the KTsP1 Central Post Computer #1 in the SM, monitoring the laptop display during ground commanding.

The FE completed the daily routine maintenance of the SM’s environment control & life support system (SOZh), including its toilet system (ASU).

After powering up the TVIS (Treadmill with Vibration Isolation & Stabilization) in the SM, the crew

completed their regular 2.5-hr. physical exercise program on the treadmill, RED (Resistive Exercise Device) and VELO bike with bungee cord load trainer.   [Valery’s daily protocol prescribes a strict four-day microcycle exercise with 1.5 hr on the treadmill and one hour on VELO plus load trainer (today: Day 4 of the first set).]

Afterwards, McArthur transferred the TVIS and RED exercise data files to the MEC (Medical Equipment Computer) for downlink, as well as the daily wristband HRM (heart rate monitor) data of the RED workout, followed by their erasure on the HRM storage medium (done six times a week).

At ~10:30am EST, the crew had their fifth regular (nominally weekly) teleconference with the Lead Flight Director at MCC-H and TsUP/Moscow via S-band/audio, with a phone patch between Houston and Moscow.

Weekly Science Update (Expedition Twelve — 6th)

Human Research Facility/Gas Analyzer System for Metabolic Analysis Physiology (HRF GASMAP):   Next week the crew will perform the GASMAP Routine Health Check, which is done to maintain the internal vacuum.  The ground team will be on console for support..

Human Research Facility/Workstation (HRF WS):  This week the Science Officer successfully completed the Workstation 2 functional checkout. The Workstation team thanked him for his efforts, “pleased to know that all operations were nominal”.

Pulmonary Function System (PFS):  Continuing.

Advanced Ultrasound (ADUM):  Continuing.

Renal Stone (RS):    Continuing.

Foot/Ground Reaction Forces during Space Flight (FOOT):   Complete for this Increment.

Space Acceleration Measurement System (SAMS):   SAMS is powered off.

Microgravity Acceleration Measurement System (MAMS):   MAMS is powered off.

Binary Colloidal Alloy Test-3 (BCAT-3):    Over the past two weeks Bill McArthur has taken photos of all of the viable samples on BCAT (samples 1-6 and 8-10).  Specialists are hopeful that revisiting samples 8-10 will find surface crystals which are difficult to see (but worth the effort).  The photos of samples 1-6 showed that more than 4 samples may have phase separated.  This is very exciting, and the team is hoping to get more pictures of BCAT sample 4 (and possibly 5 which may have shown an interface).  Seeing the interface clearly in BCAT sample 4 (and in sample 5 if an interface exists in that sample) is very important to the specialists since these are the samples nearest the critical point.

Materials ISS Experiment (MISSE):  In progress.  New MISSE-5 “suitcase” deployed and unfolded during EVA outside on the U.S. Airlock.  Temporarily turned off during the Soyuz relocation.

Dust and Aerosol Measurement Feasibility Test (DAFT):   Nothing new.

Earth Knowledge Acquired by Middle School Students (EarthKAM):   Planned for February.

Serial Network Flow Monitor (SNFM):  Complete.

Fluid Merging Viscosity Measurement (FMVM):  Nothing new.

Space Experiment Module (SEM):  Nothing new.  Experimenters and kids are working to get the next two satchels on ULF1.1.

Miscible Fluids in Microgravity (MFMG):   MFMG payload operations are finished.

Educational Payload Operations (EPO):   Complete.

Crew Earth Observations (CEO):   Through 11/13, the ground has received and reviewed a total 2,443 CEO images for this Increment.  Targets acquired this week included the Florida Coastal Everglades, LakePoopo, Moorea Coral Reef, the Muglad Basin Fans, and the Sobat Fans.  “Both your technique and the composition of your imagery are excellent.”


ISS Location NOW

Full Size/Update
Real Time ISS TrackerMore Links

Today’s CEO (Crew Earth Observations) photo targets, in the current LVLH attitude no longer limited by flight rule constraints on the use of the Lab nadir/science window, were Internal waves, E & W Florida coasts (on this pass the crew was to look well right of track for glint-enhanced features of the SE Florida coast from the Miami-West Palm Beach Area to the northern Bahamas), Georgia Coastal Ecosystems (if possible after looking right of track for sun glint, the crew was to look just left of track to carefully map the coastal islands and estuaries of Georgia.  This area [between Jacksonville, Florida and Hilton Head, South Carolina] is part of a Long Term Ecological Research [LTER] site), Red River Basin, TX (the weather should remain clear over north Texas and southern Oklahoma.  The area of interest is that portion of the Red River east of Palo Duro Canyon.  Using the long lens and documenting land use changes with a mapping pass eastward towards Lake Texoma), and Konza Prairie, Kansas (this Long Term Ecological Research [LTER] site covers most of eastern Kansas to the west and south of the Missouri River.  Because of its stony soils, it has been less agriculturally developed than other areas in the US Great Plains, and stands of original prairie grasses remain there.  As ISS crossed the Arkansas River, the crew was to begin near-nadir mapping until reaching the Kansas City area.

To date, over 177,000 of CEO images have been taken in the first five years of the ISS.

 CEO photography can be viewed and studied at the websites:

See also the website “Space Station Challenge” at:

To view the latest photos taken by the expedition 12 crew visit:

Expedition 12 Flight Crew Plans can be found at

Previous NASA ISS On-orbit Status Reports can be found here. Previous NASA Space Station Status Reports can be found here. Previous NASA Space Shuttle Processing Status Reports can be found here. A collection of all of these reports and other materials relating to Return to Flight for the Space Shuttle fleet can be found here.

Upcoming Events (all dates Eastern):

  • 12/20/05 — Progress M-54/19P undocking & reentry
  • 12/21/05 — Progress M-55/20P launch
  • 12/23/05 — Progress M-55/20P docking
  • 01/09/06 — 100 days for Expedition 12.
  • end-Jan./06 — EVA-15 (Russian; date TBD).

ISS Altitude History

Apogee height Mean AltitudePerigee height

ISS Altitude History

For more on ISS orbit and worldwide ISS naked-eye visibility dates/times, see In addition, information on International Space Station sighting opportunities can be found at on NASA’s Human Spaceflight website. The current location of the International Space Station can be found at at NASA’s Marshall Space Flight Center. Additional satellite tracking resources can be found at

SpaceRef staff editor.